Nancy Bondy's aunt has promised to deposit $1 in her account on the first day of her birthday month, $2 on the second day, $3 on the third day, and so on for 30 days. How much will this amount to over the entire month.
Answer by jim_thompson5910(35256)

This is an arithmetic sequence

So use the arithmetic series formula


S = n*(a1+an)/2

S = 30*(1+30)/2

S = 15*31

S = 465


There will be $465 in the account when the last deposit (the 30th deposit) is made.